Canada Golf Club (CGC) is considering three independent projects for July 2021 tournament. The three projects are project A, project B and project C. Given the following cash flow information, calculate the payback period for each. If CGC requires a 3-year payback before an investment can be made, which project(s) would be accepted?

Year 0 (investment) 1 2 3 4 5

Project A -2,000 -2,000 800 600 600 400

Project B -10,000 -6,000 4,000 3,000 2,000 2,000

Project C -5,000 -2,000 5,000 5,000 5,000 2,000


1
Expert's answer
2021-04-19T18:50:18-0400

Project A: "-2,000-2,000+800+600+600+400=-1,600" does not payback

Project B: "-10,000-6,000+4,000+3,000+2,000+2,000=-5,000" does not payback

Project C: "-5,000-2,000+5,000=-2,000"; "\frac{2,000}{5,000}*12=4.8" months;

payback period 2 years and 5 months

Project C will be accepted